# Bounds on Dao numbers and applications to regular local rings

### v1

@@ 4. Approach (and a conjecture) via reduction numbers
(((([17, Corollary 3.7])))) If $(R,\mathfrak{m})$ is a one-dimensional Cohen-Macaulay local ring with infinite residue field, then
@@
It is worth pointing out that this proposition is no longer true if the reduction $I$ is not required to be minimal, as exemplified in [17, Example 4.2].
@@
(((([17, Conjecture 3.8])))) If $(R,\mathfrak{m})$ is a Cohen-Macaulay local ring with infinite residue field and ${\rm dim}\,R\geq 2$, then
@@
(((([17, Theorem 3.4])))) Let $(R,\mathfrak{m})$ be a local ring with infinite residue field and ${\rm depth}\,R>0$, and let $I$ be a reduction of $\mathfrak{m}$. Then
@@
Let $(R,\mathfrak{m})$ be the local ring of a rational triple point as in [17, Example 4.1], where we highlighted that $s(\mathfrak{m})=1$. Then, by Theorem 4.3, we obtain $\mathfrak{d}_{3}(I)={\rm r}_{I}(\mathfrak{m})$ for any reduction (in particular, minimal reduction) $I$ of $\mathfrak{m}$.
@@
Since $I$ is a minimal reduction of $\mathfrak{m}$, we can apply [14, Proposition 2.4 and Theorem 2.10] to obtain $I\widetilde{\mathfrak{m}^{k}}=\widetilde{\mathfrak{m}^{k+1}}$ for all $k\geq{\rm r}_{I}(\mathfrak{m})$. Consequently,
@@
Proof. For such $R$, the ring ${\rm gr}_{\mathfrak{m}}(R)$ is Cohen-Macaulay (see [22, Theorem 2]), hence its depth is equal to ${\rm dim}\,R\geq 2>0$. In this case, as we already know, the conjecture holds true. $\square$
@@ 5. Application: Regular local rings
In this section, as applications of some of our results, we provide new characterizations of regular local rings, and describe a potential approach to the long-standing Zariski-Lipman conjecture about derivations.
@@ 5.1. Characterizations of regular local rings
Our result in this part is Theorem 5.3 below. We observe that the implication (a) $\Rightarrow$ (d) recovers [17, Corollary 3.11]. Moreover, a crucial fact here (which, as far as we know, is new) is given by (b) $\Rightarrow$ (a), i.e., $(R,\mathfrak{m})$ must be regular if $\mathfrak{m}$ is generated by a $d$-sequence (see [11], [12] for the definition of this type of sequence and its properties), which in particular solves the problem suggested in [17, Remark 3.12]. Finally, the equivalence between assertions (a) and (e) reveals the curious role played by ${\mathfrak{d}}_{3}(I)$ in regard to the theory of regular local rings, which can be re-expressed by means of equivalence to the structural assertion (f).
@@
([23, Corollary 5.2]) Let $(R,\mathfrak{m})$ be a local ring with infinite residue field. Then, $\mathfrak{m}$ is generated by a $d$-sequence if and only if $\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\mathcal{R}(\mathfrak{m})=0$.
@@
([15, Theorem 1.1]) Let $(R,\mathfrak{m})$ be a local ring with ${\rm depth}\,R>0$. If $Q$ is a parameter ideal of $R$ such that $Q^{n}$ is $\mathfrak{m}$-full for some $n\geq 1$, then $R$ is regular.
@@
(b) $\Rightarrow$ (c) According to [20, Theorem 2.1(ii)], we have ${\rm max}\{\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\mathcal{R}(\mathfrak{m}),1\}\geq s(\mathfrak{m})$. Hence, using Lemma 5.1, we obtain $s(\mathfrak{m})=1$. In order to deal with ${\rm r}_{I}(\mathfrak{m})$, consider first the case where the reduction $I$ is minimal. Applying [20, Theorem 1.3] and [21, p. 12], we derive
@@
(c) $\Rightarrow$ (d) This follows readily from the fact (proved in [17, Theorem 3.4]) that
@@
Now, by [17, Proposition 2.2], the vanishing of ${\mathfrak{d}}_{3}(I)$ forces that of ${\mathfrak{d}}_{1}(I)$ and ${\mathfrak{d}}_{2}(I)$.
@@
(e) $\Rightarrow$ (a) In the present setting, $\mathfrak{m}$ admits a reduction $Q$ which is a parameter ideal (see, e.g., [13, Exercise 8.11(ii)]). As ${\mathfrak{d}}_{3}(Q)={\mathfrak{d}}_{1}(Q)$, we obtain
@@ 5.2. Potential approach to a classical conjecture
This problem has an interesting long history, featuring in particular a strong geometric counterpart, and remains open in some cases. For details and references, see [9, Section 2] (in particular, see [9, Theorem 2.3] for a simple proof of the Zariski-Lipman conjecture in the graded case). Additionally, [16, Section 4] shows a relation between this conjecture and the $\mathfrak{m}$-full property of ideals in the (open) two-dimensional local case, which makes it somewhat natural to expect further connections.
@@
Acknowledgements. The second-named author was partially supported by CNPq (grants 406377/2021-9 and 313357/2023-4).

### v2

@@ 4. Approach (and a conjecture) via reduction numbers
(((([18, Corollary 3.7])))) If $(R,\mathfrak{m})$ is a one-dimensional Cohen-Macaulay local ring with infinite residue field, then
@@
It is worth pointing out that this proposition is no longer true if the reduction $I$ is not required to be minimal, as exemplified in [18, Example 4.2].
@@
(((([18, Conjecture 3.8])))) If $(R,\mathfrak{m})$ is a Cohen-Macaulay local ring with infinite residue field and ${\rm dim}\,R\geq 2$, then
@@
(((([18, Theorem 3.4])))) Let $(R,\mathfrak{m})$ be a local ring with infinite residue field and ${\rm depth}\,R>0$, and let $I$ be a reduction of $\mathfrak{m}$. Then
@@
Let $(R,\mathfrak{m})$ be the local ring of a rational triple point as in [18, Example 4.1], where we highlighted that $s(\mathfrak{m})=1$. Then, by Theorem 4.3, we obtain $\mathfrak{d}_{3}(I)={\rm r}_{I}(\mathfrak{m})$ for any reduction (in particular, minimal reduction) $I$ of $\mathfrak{m}$.
@@
Since $I$ is a minimal reduction of $\mathfrak{m}$, we can apply [15, Proposition 2.4 and Theorem 2.10] to obtain $I\widetilde{\mathfrak{m}^{k}}=\widetilde{\mathfrak{m}^{k+1}}$ for all $k\geq{\rm r}_{I}(\mathfrak{m})$. Consequently,
@@
Proof. For such $R$, the ring ${\rm gr}_{\mathfrak{m}}(R)$ is Cohen-Macaulay (see [23, Theorem 2]), hence its depth is equal to ${\rm dim}\,R\geq 2>0$. In this case, as we already know, the conjecture holds true. $\square$
@@ 5. Application: Regular local rings
In this section, we provide new characterizations of regular local rings and describe a potential approach to the long-standing Zariski-Lipman conjecture about derivations.
@@ 5.1. Characterizations of regular local rings
Our result in this part is Theorem 5.3 below. We observe that the implication (a) $\Rightarrow$ (d) recovers [18, Corollary 3.11]. Moreover, a crucial fact here (which, as far as we know, is new) is given by (b) $\Rightarrow$ (a), i.e., $(R,\mathfrak{m})$ must be regular if $\mathfrak{m}$ is generated by a $d$-sequence (see [12], [13]), which in particular solves the problem suggested in [18, Remark 3.12]. Finally, the equivalence between assertions (a) and (e) reveals the curious role played by ${\mathfrak{d}}_{3}(I)$ in regard to the theory of regular local rings, which can be re-expressed by means of equivalence to the structural assertion (f).
We recall, for completeness, that a sequence of elements $x_{1},\ldots,x_{m}\in R$ is said to be a $d$-sequence if no element of the sequence lies in the ideal generated by the others and, in addition, there are equalities $0:_{R}x_{1}x_{j}=0:_{R}x_{j}$ for $j=1,\ldots,m$ and $(x_{1},\ldots,x_{i}):_{R}x_{i+1}x_{j}=(x_{1},\ldots,x_{i}):_{R}x_{j}$ for $i=1,\ldots,m-1$ and $j=i+1,\ldots,m$. This holds, e.g., if $x_{1},\ldots,x_{m}$ is a regular sequence, but there are plenty of examples of $d$-sequences that are not regular.
@@
([24, Corollary 5.2]) Let $(R,\mathfrak{m})$ be a local ring with infinite residue field. Then, $\mathfrak{m}$ is generated by a $d$-sequence if and only if $\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\mathcal{R}(\mathfrak{m})=0$.
@@
([16, Theorem 1.1]) Let $(R,\mathfrak{m})$ be a local ring with ${\rm depth}\,R>0$. If $Q$ is a parameter ideal of $R$ such that $Q^{n}$ is $\mathfrak{m}$-full for some $n\geq 1$, then $R$ is regular.
@@
(b) $\Rightarrow$ (c) According to [21, Theorem 2.1(ii)], we have ${\rm max}\{\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\mathcal{R}(\mathfrak{m}),1\}\geq s(\mathfrak{m})$. Hence, using Lemma 5.1, we obtain $s(\mathfrak{m})=1$. In order to deal with ${\rm r}_{I}(\mathfrak{m})$, consider first the case where the reduction $I$ is minimal. Applying [21, Theorem 1.3] and [22, p. 12], we derive
@@
(c) $\Rightarrow$ (d) This follows readily from the fact (proved in [18, Theorem 3.4]) that
@@
Now, by [18, Proposition 2.2], the vanishing of ${\mathfrak{d}}_{3}(I)$ forces that of ${\mathfrak{d}}_{1}(I)$ and ${\mathfrak{d}}_{2}(I)$.
@@
(e) $\Rightarrow$ (a) In the present setting, $\mathfrak{m}$ admits a reduction $Q$ which is a parameter ideal (see, e.g., [14, Exercise 8.11(ii)]). As ${\mathfrak{d}}_{3}(Q)={\mathfrak{d}}_{1}(Q)$, we obtain
@@
The equivalence (a) $\Leftrightarrow$ (e) of Theorem 5.3 leaves the following natural problem.
Question 5.4.
Let $(R,\mathfrak{m},K)$ be either a local ring or a standard graded $K$-algebra, with $K$ infinite and $\mathrm{depth}\,R>0$. What kind of rings could be characterized by the property that ${\mathfrak{d}}_{3}(I)\leq 1$ for all minimal reductions $I$ of $\mathfrak{m}$?
Though quite vaguely and having no indication of proof, the authors suspect that the answer to the question should rely in the nature of the singular locus of $R$, for instance, the situation where $R$ has at most an isolated singularity at the origin (the case of rational singularities should not be ruled out either).
@@ 5.2. Potential approach to a classical conjecture
This problem has an interesting long history, featuring in particular a strong geometric counterpart, and remains open in some cases. For details and references, see [9, Section 2] (in particular, see [9, Theorem 2.3] for a simple proof of the Zariski-Lipman conjecture in the graded case). Additionally, [17, Section 4] shows a relation between this conjecture and the $\mathfrak{m}$-full property of ideals in the (open) two-dimensional local case, which makes it somewhat natural to expect further connections.
@@
Acknowledgements. The first-named author was partly supported by the Grant JDC2023-051705-I funded by MICIU/AEI/10.13039/501100011033 and by the FSE+. The second-named author was partially supported by CNPq (grants 406377/2021-9 and 313357/2023-4). The authors are grateful to the referee for a number of helpful comments and suggestions, and for pointing out an inaccuracy in an earlier proof of Theorem 3.4.